40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 1646     Entry time: Wed Jun 3 03:30:52 2009
Author: rana 
Type: Update 
Category: MOPA 
Subject: NPRO current adjust 
I increased the NPRO's current to the max allowed via EPICS before the chiller shutdown. Yesterday, I did this
again just to see the effect. It is minimal.

If we trust the LMON as a proportional readout of the NPRO power, the current increase from 2.3 to 2.47 A gave us
a power boost from 525 to 585 mW (a factor of 1.11). The corresponding change in MOPA output is 2.4 to 2.5 W
( a factor of 1.04).

Therefore, I conclude that the amplifier's pump has degraded so much that it is partially saturating on the NPRO
side. So the intensity noise from NPRO should also be suppressed by a similar factor.

We should plan to replace this old MOPA with a 2 W Innolight NPRO and give the NPRO from this MOPA back to the
bridge labs. We can probably get Eric G to buy half of our new NPRO as a trade in credit.
Attachment 1: Untitled.png  171 kB  | Hide | Hide all
Untitled.png
ELOG V3.1.3-